1. Express the displacement of velocity and acceleration as a function of time


2. Express velocity and acceleration as a function of time


3. A mass of 100g is connected is a light spring of force constant 10Nm and is free to oxillate on a horizontal frictionless surface. If the mass is displaced 8cm from equilibrium and release from the rest. Find: i. The period of it's motion ii. The maximum speed of the mass iii. The maximum acceleration of the mass


1
Expert's answer
2021-05-06T17:22:04-0400

Solution :-

Displacement "x=asin(wt+\phi)"

(a) Velocity in displace ment form

"v=\frac{dx}{dt}=aw"

Acceleration in displacement form


"a=\frac{d^2x}{dt^2}=w^2x"

(b) Acceleration in term of velocity

"a=\frac{dv }{dt}=w^2x"


(c) we know that

Kx =mg

We know that

"T=2\pi \sqrt\frac
{m}{k}"

Put value

"T" "=2\pi \sqrt\frac
{m}{k}=2\pi \sqrt\frac
{100\times10^{-3}}{10}=0.628sec"

T=0.628sec

Maximum speed 0.8m/sec

Maximum speed(v)="\omega x=2\pi n(x)"

"v=\frac{2\pi}{T}x"

"v=\frac{2\pi}{0.628}\times.08=0.8m/sec"

Put values

V=0.8m/sec

Maximum acceleration

"a=\omega^2x"

Put value "a=(2\pi n)^2x"

Put value

"a=(\frac{2π}{T})^2x"

"a=(\frac{2\times3.14}{0.628})^2\times0.08"

"a=8m\/sec^2"

"a=9.80m\/sec^2"